# HLE-Verified: A Systematic Verification and Structured Revision of Humanity’s Last Exam

### v2

@@ A.3 LLM Judge Prompt in Stage II
A.4 Quality control and decision principles
@@
A.5 Case Study
@@
A.6 Component-wise Defect Distribution Across Subjects
@@
A.7 Cross-subject proportional differences.
@@
Mathematics. Answer-level defects are almost entirely Incorrect Answer (over 95%). At the rationale level, Missing Prerequisite and structural incompleteness dominate. Problem-level defects remain comparatively limited. This pattern indicates that mathematical items are generally well-posed but frequently suffer from incorrect answer keys and incomplete derivations.
Biology/Medicine. Biology/Medicine exhibits a similarly high dominance of incorrect answers, while rationale-level incompleteness is particularly pronounced. Compared to mathematics, format-level issues are less central, and implicit domain assumptions play a larger role.
Chemistry. Chemistry presents a more balanced distribution. While incorrect answers remain the leading answer-level defect, both problem- and rationale-level format semantic errors occupy a larger share than in other domains, reflecting sensitivity to symbolic and notational precision.
Physics. Physics shows fewer total defects and a flatter proportional structure. Although incomplete rationales remain the leading issue, answer-level dominance is less extreme than in mathematics and biomedicine, suggesting comparatively stronger answer-key stability.
Computer Science. Computer Science exhibits a distinctive pattern in which format semantic errors dominate both problem- and rationale-level defects. This reflects the domain’s reliance on representation conventions, indexing standards, and symbolic alignment.

### v3

@@ A.1.2 LLM Judge Prompt in Stage II
A.2 Quality control and decision principles
@@
Expert recruitment and profile.
The verification and revision process was conducted by domain experts recruited through two independent supplier teams and supported by internal adjudication specialists. The participating experts are predominantly Master’s- and Ph.D.-level researchers from research-intensive universities and institutes, with formal academic training in mathematics, physics, chemistry, biomedicine, and computer science. Their academic backgrounds align directly with the disciplinary scope of HLE-Verified, enabling subject-matter–grounded validation and revision across all covered domains. Independent acceptance reviewers with doctoral-level expertise further examined complex or contentious cases.
A.3 Case Study
@@
A.4 Component-wise Defect Distribution Across Subjects
@@
A.5 Cross-subject proportional differences.
@@
Mathematics. Answer-level defects are overwhelmingly dominated by Incorrect Answer (89.4%). At the rationale level, type 3 (40.1%) is the leading defect, reflecting structural incompleteness in derivations. Problem-level defects are more moderate in scale, with type 5 (40.0%) most frequent. This pattern suggests that mathematical items are generally conceptually stable but vulnerable to answer-key inaccuracies and incomplete reasoning articulation.
Physics. Physics exhibits relatively small defect counts overall. Answer-level errors remain dominated by type 1 (80.0%). Problem-level defects are led by type 1 (57.1%), while rationale defects show a flatter distribution with type 3 accounting for 25.0%. Compared to other domains, physics demonstrates comparatively less concentration in a single rationale category.
Chemistry. Chemistry shows a more balanced defect structure. Although incorrect answers (71.4%) remain the leading answer-level issue, both problem- (56.5%) and rationale-level (40.0%) defects are dominated by format semantic errors. This reflects the domain’s sensitivity to symbolic precision and representational alignment.
Biology/Medicine. Biology/Medicine displays the highest dominance of incorrect answers (97.2%). Rationale-level defects are primarily type 3 (45.0%), indicating incomplete explanatory structure. Problem-level defects are more evenly distributed, with type 2 (33.3%) emerging as the most frequent category.
Computer Science. Computer Science exhibits the most pronounced representation sensitivity. Format semantic errors (type 5) dominate problem-level defects (94.6%), while type 10 dominates rationale-level defects (63.3%). Answer-level incorrectness (78.1%) remains substantial but less extreme than in mathematics or biomedicine.